HUBBLE SPACE TELESCOPE DAILY REPORT #2482 PERIOD COVERED: 0000Z (UTC) 10/19/99 - 0000Z (UTC) 10/20/99 Daily Status Report as of 293/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed Two Sets of WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.3 Completed Three Sets of STIS/CCD/MA1 8204 (Calibration of the Wind Momentum-Luminosity Relationship in Cyg OB2) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations of the stellar winds around the O/B class stars Cyg OB2-8A, Cyg OB2-10 and Cyg OB2-2. A precise calibration of the Wind Momentum-Luminosity Relationship using objects of the solar neighborhood is crucial for future applications of the method to determine extragalactic distances out to the Virgo and Fornax galactic clusters. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. The observations were completed as scheduled, and no problems were reported. 1.5 Completed Four Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was to be used to make a CCD bias calibration observation. As described in 2.1 and HSTAR 7322, the acquisition for this proposal failed, the take data flag remained down, and two observations were affected. 1.7 Completed WF/PC-2 8183 (Morphology of the Most Massive Galaxies in the Early Universe) The WF/PC-2 was used to make observations of the massive radio galaxy MRC2224-27 in order to study its morphology. High redshift radio galaxies are fundamentally important probes of galaxy and cluster formation. As described in 2.1 and HSTAR 7322, the acquisition for the first portion of this proposal failed, the take data flag remained down, and two observations were affected. The successful re-acquisition occurred midway through this proposal. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.8 Completed STIS/MA2 8426 (Cycle 8 MAMA Dark Measurements) The Space Telescope Imaging Spectrograph (MA2) was used to make a routine Cycle-8 MAMA dark noise calibration measurement. This proposal will provide the primary means of checking on the health of the MAMA detectors. This is done through frequent monitoring of the background count rate. The observations were completed as planned, and no problems were reported. 1.9 Completed Three Sets of STIS/CCD 8549 (POMS Test Proposal: STIS non-scripted parallel proposal continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude observations as part of a non-scripted POMS Test Proposal. The observations were completed as scheduled, and no problems were reported. 1.10 Completed WF/PC-2 8151 (NGC 1866: A Critical Test of Stellar Evolution for Intermediate Mass Stars) The WF/PC-2 was used to make observations of LMC cluster NGC-1866 in order to help study the stellar evolution for intermediate mass stars. The cluster NGC 1866 contains at least 23 Cepheid variable stars. The observations were executed as scheduled, and no anomalies were noted. 1.11 Completed STIS/CCD 8545 (POMS Test Proposal: STIS non-scripted parallel proposal continuation VI) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.12 Completed Two Sets of STIS/CCD 8037 (STIS Coronographic Imaging of Circumstellar Discs) The Space Telescope Imaging Spectrograph (CCD) was used to make coronographic observations of the circumstellar protoplanetary disk around the star CQ-Tauri. The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed WF/PC-2 8216 (T Tauri Star Snapshot Survey II. Completing the CTTs Sample) The WF/PC-2 was used to make observations of the T-Tauri class star known as L1642-2 as part of Snapshot Survey II. This database will be of fundamental importance to understanding the nature of planet-forming environments and their frequency in the galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.14 Completed WF/PC-2 8332 (A Snapshot Survey of Symbiotic Stars) The WF/PC-2 was used to make observations of the star V1329-Cyg as part of a snapshot survey of symbiotic stars. A fuller understanding of the symbiotic star phenomenon is important to several branches of present day astrophysics including mass loss, accretion onto compact objects, and photoionization and radiative transfer in gaseous nebulae. The observations were completed as planned, and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 10 Successful: 9 The acquisition scheduled for 292/103000Z failed to RGA control when the scan step limit was exceeded for FGS-1. HSTAR 7322 was written. The following re-acquisition at 292/115720Z was successful. The proposals described in 1.6 and 1.7 were affected by this event. Scheduled Re-acquisitions: 5 Successful: 5 2.2 FHST Updates: Scheduled: 20 Successful: 20 2.3 Operations Notes: The SSR EDAC error counter was cleared twice per ROP SR-1. The NICMOS FPA temperature sensor for detector 2 was enabled at 292/1658Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH